# DGE-YOLO: Dual-Branch Gathering and Attention for Accurate UAV Object Detection

### v1

@@ IV-D Comparison with State-of-the-art Methods
Figure 4: Visualization results. Compared our method with results of baseline in two unimodal. The red detection box represents the car class, pink represents the truck class, purple represents the freight car class, and orange represents the bus class. To better distinguish between unimodal and multimodal detection bounding boxes, different anchor box colors are used. The red dashed boxes represent misdetected targets, the yellow dashed boxes represent missed targets, and the green dashed boxes represent correctly detected targets.
TABLE II: Ablation Studies.IR and RGB represent infrared single mode and visible light single mode respectively, and when selected at the same time, they represent the fusion of the two modes.
@@ IV-E Visualization
In this section, we validate the effectiveness of the model through visual demonstrations. Since image fusion has not been applied, we chose the infrared modality, which contains more feature information, to represent our model for performance display. As shown in Fig.4, under dark conditions, many targets are difficult to detect in the visible light modality, resulting in a significant loss of targets. From the detection results, it can be observed that the original model has different advantages when detecting the same target in the visible light and infrared modalities. Overall, the proposed model not only addresses the issue of missed detections effectively but also improves the recognition of false positives.
IV-F Ablation Studies and Analysis
In this section, we conduct ablation experiments to validate the effectiveness of each module in the model. The experimental results are presented in TABLEII. In what follows, This section will comprehensively analyze the proposed DGE-YOLO method from three aspects to explore the logic behind its superior performance. (1) Role of dual-branch network. The detection accuracy of the bimodal approach is higher than both the infrared and visible light single-modality detection accuracies. This indicates that the infrared modality contains more feature information than the visible light modality, and that the bimodal approach has the ability to fuse a wider range of features. After the application of the dual branch network, the Gather-and-Distribute module and EMA module can more effectively improve the performance of the model. The performance of the dual channel model is higher than that of the single modal model, and the experimental results are shown in Fig.5.
Figure 5: Introduce comparisons of different modules in different modalities.
Figure 6: Introduce comparisons of different modules in different modalities.
(2) Impact of EMA. while the EMA module performs better in improving mAP50:95, As shown in Fig.5. MAP50:95 provides a more comprehensive evaluation by considering a range of IoU thresholds from 0.5 to 0.95. This indicates that the model’s performance has improved across varying degrees of object overlap. But also because the module performs better in dealing with complex fusion Feature, it also performs generally when single mode is introduced. (3) Influence of Gather-and-Distribute module. The experimental results show that the Gather-and-Distribute module is more effective in improving mAP50, As shown in Fig.6. this module effectively improves the overall performance of the model. However, since the modified module is used to process the feature information after feature fusion, it has a general effect when introduced in the case of single mode.
V Conclusion
This paper proposes an improved model DGE-YOLO, based on YOLOv8n, for multimodal drone detection tasks. The original model is modified from a single-modality to a dual-input model, making it suitable for handling both infrared and visible light tasks. By improving the network structure, the model effectively combines the advantages of different modalities, achieving the best performance in drone detection tasks while maintaining lower computational complexity compared to most other models. Through extensive experiments and data analysis, The map50 of the model reached 84.7% with 7.1 parameters and 15.2 Flops, the proposed model shows clear advantages over the latest single-modal and multimodal models, demonstrating its advanced capabilities. However, the method used in this paper for handling different modality feature layers is relatively simple. Future work can focus on this issue as a research direction, exploring how to better integrate features from different modalities to further enhance the model’s ability to process multimodal data.

### v2

@@ 3 Experiment and Analysis
Fig. 4: Visualization results. Compared our method with results of baseline in two unimodal. The red detection box represents the car class, pink represents the truck class, purple represents the freight car class, and orange represents the bus class. The red dashed boxes represent misdetected targets, the yellow dashed boxes represent missed targets, and the green dashed boxes represent correctly detected targets.
3.1 Datasets
Drone Vehicle. This paper evaluates our method on the Drone Vehicle dataset, The dataset includes five object categories: car, truck, bus, freight car, and van. Following the training approach in the original paper, we used 17,990 image pairs for training and 1,469 image pairs for validation.
3.2 Comparison with State-of-the-art Methods
The experimental results are shown in Table1. Among the single-modal detection models, the detection performance under the infrared modality is significantly higher than that under the visible light modality. We selected the YOLOv8 and subsequent versions to validate the effectiveness of our model. Compared to previous detection models, our model exhibits a significant improvement in accuracy. Since the parameters and flops of the single-stage detection model are much smaller than those of the multimodal fusion model, no comparison is made in the experiment. Our model has shown substantial improvements in categories with previously lower accuracy, achieving the highest mAP50, and a significant boost in the mAP50:95 metric.This paper proposed DGE-YOLO method outperforms many state-of-the-art single-modal and multimodal object detection methods.
Table 2: Ablation Studies.
@@ 3.2 Comparison with State-of-the-art Methods
Visualization. In this section, we validate the effectiveness of the model through visual demonstrations. Since image fusion has not been applied, we chose the infrared modality, which contains more feature information, to represent our model for performance display. As shown in Fig.4, under dark conditions, many targets are difficult to detect in the visible light modality, resulting in a significant loss of targets. From the detection results, it can be observed that the original model has different advantages when detecting the same target in the visible light and infrared modalities. Overall, the proposed model not only addresses the issue of missed detections effectively but also improves the recognition of false positives. Ablation Studies. In this section, we conduct ablation experiments to validate the effectiveness of each module in the model. The experimental results are presented in TABLE2. The detection accuracy of the bimodal approach is higher than both the infrared and visible light single-modality detection accuracies, This indicates that the infrared modality contains more feature information than the visible light modality.while the EMA module performs better in improving mAP50:95, As shown in Fig.5(a). Because the module performs better in dealing with complex fusion Feature, it also performs generally when single mode is introduced. The experimental results show that the Gather-and-Distribute module is more effective in improving mAP50, this module effectively improves the overall performance of the model, As shown in Fig.5(b).
Fig. 5: Introduce comparisons of different modules in different modalities.
4 Conclusion
We presented DGE YOLO, a multi modal UAV detector that combines a dual branch backbone for infrared and visible inputs, an Efficient Multi scale Attention module to strengthen cross scale semantics, and a Gather and Distribute neck that reduces aggregation loss. Experiments on the Drone Vehicle dataset show consistent gains over strong baselines and prior state of the art, particularly for small objects in complex scenes. The framework is compatible with existing YOLO variants and offers a practical solution for robust and efficient aerial perception.